In Hotelling’s town, if all firms are required to charge the same fixed price, describe the equilibrium location of three firms. Explain your answer and then describe the equilibrium for four firms.
Answer #1

ANSWER:-

  • In Hotelling's town if all firms are required to charge the equivalent fixed price and the purchasers are spread uniformly then the equilibrium location of the firms ought to be should that they should locate to find with the end goal that they could serve equivalent number of purchasers.
  • With three firms in the market they should locate to find so they could serve 1/3 of clients each. Assuming that the client will arrive at the firm closest to it.
  • The first firm should locate to situate at 1/6, second should locate to situate at 1/6+1/3=3/6=1/2 and the third firm should locate to situate at 1/2 +1/3=5/6.
  • In the comparable manner, If there are four firms they should serve 1/4 clients each. First firm should loacte to situate at 1/8, second at 1/8+1/4=3/8, third at 3/8+1/4=5/8 and fourth firm should locate to situate at 5/8+1/4=7/8.

PLEASE UPVOTE.